# Survey of Natural Language Processing for Education: Taxonomy, Systematic Review, and Future Trends

### v3

@@ Abstract
Natural Language Processing (NLP) aims to analyze text or speech via techniques in the computer science field. It serves the applications in domains of healthcare, commerce, education and so on. Particularly, NLP has been widely applied to the education domain and its applications have enormous potential to help teaching and learning. In this survey, we review recent advances in NLP with the focus on solving problems relevant to the education domain. In detail, we begin with introducing the related background and the real-world scenarios in education where NLP techniques could contribute. Then, we present a taxonomy of NLP in the education domain and highlight typical NLP applications including question answering, question construction, automated assessment, and error correction. Next, we illustrate the task definition, challenges, and corresponding cutting-edge techniques based on the above taxonomy. In particular, LLM-involved methods are included for discussion due to the wide usage of LLMs in diverse NLP applications. After that, we showcase some off-the-shelf demonstrations in this domain. At last, we conclude with six promising directions for future research, including more datasets in education domain, controllable usage of LLMs, intervention of difficulty-level control, interpretable educational NLP, methods with adaptive learning, and integrated systems for education. We organize all relevant datasets and papers in the open-available Github Link for better review https://github.com/LiXinyuan1015/NLP-for-Education.
@@ 1 Introduction
FIGURE 1: We demonstrate the education scenarios in real-world and how NLP applications contribute to these scenarios.
@@
Natural Language Processing (NLP) refers to the branch of computer science and artificial intelligence, that empowers computers the ability to understand text and spoken words as the way human beings can. It serves real-world scenarios. For example, NLP can help in healthcare applications. A wide range of NLP techniques are being utilized to anticipate the treatment process of the patients. Question answering [1, 2] and dialogue generation [3] help conduct medical diagnosis, administrative assistants for appointments and billing. Besides, NLP can help in commerce applications. There is a growing trend of incorporating NLP techniques into analytics and customer management platforms to help learn more about the relationship between the consumers and items via sentiment analysis [4].
@@
The widespread collaboration of new technologies into the education domain revolutionize education in the future. In the education domain, NLP techniques can help education in the following ways:
@@
  • NLP technologies provide automatic key information extraction which facilitates to highlight the key points. In the cases where teachers and students read passages, tutorials, and textbooks. The key points and knowledge should be annotated or noticed from unstructured text. Automatic key information extraction can assistant to identify the key notes which help them understand and comprehend during teaching and learning.
@@
  • NLP technologies offer intelligent tutoring and recommendation. For example, when students are learning a concept, a system can help recommend relevant concepts, practices, and reading materials such that the students can find proper reading materials as references, which enables active learning and self-inspection for students.
@@
  • NLP technologies help automatic problem solving. For example, in math or science courses, students may have extensive questions towards the exercises after the classes. To avoid repeat and overloading questions for teachers, NLP technologies can conduct reasoning and answer the questions automatically. Moreover, they can give explanations to the complex questions and intermediate solutions for the problems.
@@
  • NLP technologies facilitate automatic smart content generation. For example, teachers have trouble in preparing the teaching materials from a rush of information on the Internet or textbooks. With the help of retrieval technologies, we can create a wide variety of digital learning materials, such as retrieving reading passages for students in a reading course. Similarly, given a passage in the textbook as the context and a concept as the correct answer, NLP techniques can quickly construct a multiple-choice question associated with the test question and the distractors. This could facilitate the teachers to make quick quizzes for classes.
@@
  • NLP techniques enables personalized instruction. Students may come from different levels, learn different subjects and have difficulties in different concepts. In such a case, it is vital to track the learning trajectory and provide personalized instruction to the students. Various NLP models trained on different annotated data collected from different scenarios could contribute to the personalized education.
@@
  • NLP technologies achieve AI-supported assessment and correction. For example, it is time-consuming for teachers to grade the essays and codes from students in linguistics or coding classes, by analyzing the textual contents of the assignments, NLP applications can figure out the errors in the assignment and return the corresponding correction to the students.
We display the education scenarios in real-world, relevant NLP techniques, and the potential contributions of them in Figure 1. These techniques are able to analyze students performance, allowing for more flexible learning, thus freeing up valuable time for teachers.
@@
We notice there are a few surveys related to NLP for education. However, a line of surveys focus on a certain task. Specifically, Lan et al. [5] summarize the general advancement for knowledge base question answering tasks but the domain is not restricted to education, which maybe inapplicable for educational question answering. Bryant et al. [6] summarize the cutting-edge techniques of grammatical error correction, which is a typical tasks of educational NLP. Zhang et al. [7] carefully view the technical progress of math word problem solving tasks. Messer et al. [8] systematic review the task of code programming. However, the above surveys fail to identify the significant role and distinction of these tasks applied in education domain. Another line of surveys summarize in the shallow perspectives of concepts and definition [9, 10, 11], which are in a high-level and do not shed any light on the techniques and applications. Such that they cannot serve as the good references for researchers in NLP and computer science communities.
@@
To bridge the gap, this survey provides a comprehensive overview of NLP for education from a technical point of view. (1) New taxonomy: we present a novel taxonomy for educational NLP and identify representative NLP tasks that can be applied for learning and comprehension as well as writing and assessment in education scenarios, and then discuss them in detail; (2) Comprehensive review: we collect recent papers published in well-recognized computer science venues and review their methods for solving the challenges of corresponding tasks. We further include LLM-based methods due to the wide usage in NLP applications; (3) Rich resources: we discuss the publicly-available datasets and demonstrations with their various attributes, which makes it easy for readers to get started for their researches.
@@
FIGURE 2: An overview of NLP for education. The left section displays the systematic frame of relevant NLP techniques, including data source, processing and analysis techniques, and applications. The right section demonstrates the outcomes of diverse techniques.
@@ 2 Taxonomy of NLP in Education Domain
Systems built on the basis of NLP applications and equipped with NLP techniques could perform as empowered AI tutoring, which enhances the teaching and learning experience of students, teachers, parents, or tutors. With the systems, students are able to get access to private tutoring and extra support outside the classroom, which avoids increasing the workload of their teachers.
We first demonstrate an overview of NLP for education in Figure 2, where the systematic frame of relevant NLP techniques and their corresponding outcomes are displayed. The data sources are significant as the fundamental supports for subsequent analyses and applications. Generally, we could obtain educational data from textbooks, syllabus, exams, and curricula materials, which may contain behavioral data of students, exercise questions, and guidance signals of error correction and so on. From the perspective of teaching subjects, the data resources cover subjects like chemistry, math, linguistics, computer science and so on, which could support applications in different subjects. Given the raw data resources, we could conduct processing and analysis (e.g., knowledge graph construction [12], information extraction [13], information retrieval [14], knowledge alignment [15]) to extract the key information from massive data and organize them in structured knowledge graphs (e.g., EduKG [^1] https://edukg.cn/team), databases (e.g., WordNet [^2] https://wordnet.princeton.edu/documentation) and lists, which could be applied to various downstream applications. Upon the well-organized data, NLP tasks are applied (e.g., question answering, question construction, automated assessment, and error correction) to form high-level applications to serve teachers and students in real-world scenarios. Grammarly [^3] https://app.grammarly.com, Gauthmath [^4] https://www.gauthmath.com, CodeSubmit [^5] https://codesubmit.io, AutoQG  [^6] https://autoqg.net, and EduChat [^7] https://www.educhat.top/ are possible educational products and demonstrations developed based on these NLP applications.
This survey mainly focuses on the application-level. We highlight the most representative NLP applications with the taxonomy shown in Figure 3. Considering that NLP techniques play significant roles in different phases of education, on the one hand, NLP could serve as a teaching assistant for learning and comprehension. When students learn the curriculum, the following NLP tasks could simplify the difficulties of learning.
  • Question Answering (QA), as a typical application, could not only provide tutorials to the homework for students but also save time for overloaded educators by providing feedback on questions. Based on the different goals of question answering, textbook question answering and math word problem solving are presented as the two typical subcategories of question answering. These two tasks both aim to generate answers from the given contexts automatically. They can be widespread in the education of science subjects.
  • Question Construction (QC) is to generate questions as well as the distractors given the correct answers, which helps the teachers create a wide range of test questions using pre-existing learning resources. Generally, a good test question may consist of a question and multiple candidate answers. Therefore, question generation and distractor construction are two key sub-tasks to be reviewed.
On the other hand, NLP improves writing and assessment for students. When given constructive criticism and writing prompts, students are likely to go deep into the topic. However, teachers sometimes have difficulties responding thoughtfully to students due to time limitation. But they will benefit from the following tasks.
@@
  • Error Correction (EC) for language are useful to improve language learning via explicating the correct edits. Still, taking language learning as an example, grammatical error correction and code error correction are canonical sub-tasks for EC. The methods are not simply required to provide error-free outputs based on the ill-posed inputs, more detailed instructions should be provided to shed light on the reasons or knowledge points to the errors.
@@
Please be aware that there are some other NLP tasks that are related to education. For example, tasks like text rewriting [16] and controllable text generation [17] can also help teachers to prepare suitable teaching materials. In this survey, we mainly focus on the highlighted tasks, which are widely explored NLP tasks in education domain. Our discussion is carried out based on the above taxonomy.
{forest}
for tree= forked edges, grow’=0, draw, rounded corners, node options=align=left, calign=edge midpoint, , [NLP Applications for Education, text width=2.7cm, for tree=fill=white!20 [Textbook Question Answering, text width=1.8cm, for tree=fill=white!20 [Comprehend the multi-modal contexts and complex questions, text width=2.8cm, for tree=fill=white!20 [ Propose deep learning-based methods combining RNN and CNN [18, 19]; spatial attention [20, 21, 22, 23]; compositional strategies [24]; bilinear pooling schemes [25, 26]., text width=6.5cm, node options=align=left ] ] [Supplement domain-specific knowledge, text width=2.8cm, for tree=fill=white!20 [ Pre-train models with domain knowledge [27, 28]; diagrams parsing enhancement [29, 30, 31]., text width=6.5cm, node options=align=left ] ] [Strengthen zero-shot capabilities, text width=2.8cm, for tree=fill=white!20 [ Collaborate LLMs with textbooks [1, 32, 33, 34]; develop LLMs-based multi-agents [35, 2]., text width=6.5cm, node options=align=left ] ] ] [Math Word Problem Solving, text width=1.8cm, for tree=fill=white!20 [Understand the gap between natural language and abstract expression, text width=2.8cm, for tree=fill=white!20 [Propose seq2seq [36, 37, 38, 39] framework enriched with structured encoders and decoders [40, 41]; reinforcement learning [37, 42]; knowledge distillation [43]., text width=6.5cm, node options=align=left ] ] [Solve MWPs under zero/few-shot setting, text width=2.8cm, for tree=fill=white!20 [ Propose LLMs-enhanced MWP solvers [44, 45, 46, 47, 35]; Develop LLMs-based multi-agents [48]., text width=6.5cm, node options=align=left ] ] ] [Question Generation, text width=1.8cm, for tree=fill=white!20 [Identify the key statement from the context and generate the corresponding question, text width=2.8cm, for tree=fill=white!20 [ Conduct rule-matching [49]; develop deep learning-based methods with attention [50]; lingtuistic features [51, 52, 53]; multi-modal fusion [54]; multi-task learning [55]; reinforcement learning [56]., text width=6.5cm, node options=align=left ] ] [Generate more flexible questions for adaptive education, text width=2.8cm, for tree=fill=white!20 [ Propose difficulty-control QG methods [57, 58, 59]; syllabus-aligned methods [60, 61]; knowledge-tracking QG methods [62, 63]., text width=6.5cm, node options=align=left ] ] ] [Distractor Generation, text width=1.8cm, for tree=fill=white!20 [Design context-consistent and deceptive false answers, text width=2.8cm, for tree=fill=white!20 [ Propose rule-based methods [64]; feature-based methods [65]; generation-based methods [66, 67]; ranking-based methods [68, 69, 70]., text width=6.5cm, node options=align=left ] ] ] [Automated Essay Scoring, text width=1.8cm, for tree=fill=white!20 [Model different facets of the essay, text width=2.8cm, for tree=fill=white!20 [ Design various features (e.g. length-based features [71]; lexical features [72, 71]; prompt-relevant features [73]; readability features [71]; syntactic features [71, 74]) and solve via regression and ranking [75] with enhanced training procedure [76]., text width=6.5cm, node options=align=left ] ] [Sparse annotation of essays, text width=2.8cm, for tree=fill=white!20 [ Decouple the domain-dependent and domain-independent features and aggregate heuristic supervision signals [77, 78, 79]; design enriched annotations [80, 81]., text width=6.5cm, node options=align=left ] ] ] [Automated Code Scoring, text width=1.8cm, for tree=fill=white!20 [Understanding of the long-dependency of code, text width=2.8cm, for tree=fill=white!20 [ Conduct dynamic analysis to assess the execution of the code [82, 83, 84, 85]; static analysis to predict scores with the designed features [86, 87, 88, 89]; hybrid methods combine both [90]., text width=6.5cm, node options=align=left ] ] ] [Language Error Correction, text width=1.8cm, for tree=fill=white!20 [Identify and correct grammar errors accurately, text width=2.8cm, for tree=fill=white!20 [Design manual rules and classification-based models [91, 92]; neural network models like seq2seq models [93, 94, 95, 96, 97] and seq2edit models [98, 99, 100, 101]., text width=6.5cm, node options=align=left ] ] ] [Code Error Correction, text width=1.8cm, for tree=fill=white!20 [Identify and correct errors in the long-distance dependency of the code, text width=2.8cm, for tree=fill=white!20 [Develop language modeling [102, 103]; methods with abstract syntax tree [104, 105]; denoising strategies [106]; use identifier tagging and comment information [107]; grammar constraints [108]., text width=6.5cm, node options=align=left ] ] ] ]

### v4

@@ Abstract
Natural Language Processing (NLP) aims to analyze text or speech via techniques in the computer science field. It serves applications in the domains of healthcare, commerce, education, and so on. Particularly, NLP has been widely applied to the education domain and its applications have enormous potential to help teaching and learning. In this survey, we review recent advances in NLP with a focus on solving problems relevant to the education domain. In detail, we begin with introducing the related background and the real-world scenarios in education to which NLP techniques could contribute. Then, we present a taxonomy of NLP in the education domain and highlight typical NLP applications including question answering, question construction, automated assessment, and error correction. Next, we illustrate the task definition, challenges, and corresponding cutting-edge techniques based on the above taxonomy. In particular, LLM-involved methods are included for discussion due to the wide usage of LLMs in diverse NLP applications. After that, we showcase some off-the-shelf demonstrations in this domain, which are designed for educators or researchers. At last, we conclude with five promising directions for future research, including generalization over subjects and languages, deployed LLM-based systems for education, adaptive learning for teaching and learning, interpretability for education, and ethical consideration of NLP techniques. We organize all relevant datasets and papers in the open-available Github Link for better review https://github.com/LiXinyuan1015/NLP-for-Education.
@@ 1 Introduction
Natural Language Processing (NLP) refers to the branch of computer science and artificial intelligence, that empowers computers the ability to understand text and spoken words the way human beings can. It serves real-world scenarios. For example, NLP can help in healthcare applications. A wide range of NLP techniques are being utilized to anticipate the treatment process of the patients. Question answering [1, 2] and dialogue generation [3] help conduct medical diagnosis, administrative assistants for appointments, and billing. Similarly, NLP can help in commerce applications. There is a growing trend of incorporating NLP techniques into analytics and customer management platforms to help learn more about the relationship between the consumers and items via sentiment analysis [4].
@@
In the education domain, the widespread collaboration of new technologies into the education domain revolutionizes education in the future. NLP techniques can help education in the following ways:
@@
  • NLP technologies provide automatic key information extraction from unstructured text which facilitates highlighting the key points when teachers and students read passages, tutorials, and textbooks.
@@
  • NLP technologies offer intelligent tutoring and recommendations. When students are learning a concept, a system can help recommend relevant concepts, practices, and reading materials such that the students can find proper reading materials as references, which enables active learning and self-inspection for students.
@@
  • NLP technologies help automatic problem-solving. For example, in math or science courses, students may have extensive questions about exercises after the classes. To avoid repeated and overwhelming questions for teachers, NLP technologies can conduct reasoning and answer the questions automatically.
@@
  • NLP technologies facilitate automatic smart content generation when teachers have trouble in preparing the teaching materials from a rush of information on the Internet or textbooks. Additionally, given a passage in the textbook as the context and a concept as the correct answer, NLP techniques can quickly construct a multiple-choice question associated with the test question and the distractors. This could facilitate the teachers to make quick quizzes for classes.
@@
  • NLP techniques enable personalized instruction. Students may have different background and it is vital to track the learning trajectory and provide personalized instruction to the students. Various NLP models trained on different annotated data collected from different scenarios could contribute to personalized education.
@@
  • NLP technologies enable AI-supported assessment and correction. For example, it is time-consuming for teachers to grade the essays and codes from students in linguistics or coding classes, by analyzing the textual contents of the assignments, NLP applications can figure out the errors in the assignment and return the corresponding correction to the students.
The above techniques allow for more flexible and effective learning for students, thus freeing up valuable time for teachers.
@@
We notice there are a few surveys related to NLP for education. However, a line of surveys focuses on a certain task. Specifically, Lan et al. [5] summarize the general advancement for knowledge base question answering tasks but the domain is not restricted to education. Bryant et al. [6] summarize the cutting-edge techniques of grammatical error correction, which is a typical task of educational NLP. Zhang et al. [7] carefully review the technical progress of math word problem solving tasks. Messer et al. [8] provide a systematic review of the task of code programming. However, the above surveys focus solely on a specific task and do not identify the significant distinction of these tasks applied in the education domain. Recently, we found there are some surveys conducting a specific discussion on large language models for education [9, 10], which are advanced but cover only partial NLP techniques.
@@
To provide a comprehensive overview of NLP for education from a technical point of view: (1) New taxonomy: we present a novel taxonomy for educational NLP and identify representative NLP tasks that can be applied for learning and comprehension as well as writing and assessment in education scenarios, and then discuss them in detail; (2) Comprehensive review: we collect recent papers published in well-recognized computer science venues and review their methods for solving the challenges of corresponding tasks. We further include LLM-based methods due to their wide usage in NLP applications; (3) Rich resources: we discuss the publicly-available datasets and demonstrations with their various attributes, which makes it easy for readers to get started for their research.
@@
FIGURE 1: An overview of NLP for education. The left section displays the systematic frame of relevant NLP techniques, including data sources, processing and analysis techniques, and applications. The right section demonstrates the outcomes of diverse techniques.
NLP Applications for Education Textbook Question Answering The subject-specific QA data is limited for training Fine-tune the contextual knowledge as language modeling [11, 12] and augment data from different subjects [13]. The contexts may be presented in diverse formats Incorporate the pre-processing steps of diagrams/tables/graphs as off-the-shelf toolkits [14, 15]. Zero-shot issues frequently occur Collaborate with LLMs as the agent in TQA tasks and integrate the image-to-caption modules and pre-processing toolkits [16, 17, 18]. Math Word Problem Solving Some MWPs are complex in logic Formulate the math expression as an abstract tree and re-organize the textual contexts with structures [19, 20, 21, 22, 23, 24]. There is a lack of annotated expressions in the MWPs Learn the MWPs solving with weak supervision [19, 25]; Distill the knowledge from a large pre-trained generic model [26]. Zero/few-shot complex MWPs in real-world scenarios Force LLMs to think step by step via CoT [27, 28, 29, 30, 31, 32]; Treat LLMs as agent to call toolkits for processing contexts in multi-modalities [33]. Question Generation The generated questions ignore the difficulty stratification Setup a QA model as the discriminator [34]; Propose a feature of difficulty levels for learning based on the number of inference steps/ item response theory/ Bloom’s Taxonomy [35, 36, 37]; Prompt LLMs to indicate difficulty levels [38, 39]. The generated questions show weak relevance to the syllabus Setup a binary model as the discriminator to judge if it is related to certain keywords or concepts[40, 41, 42] which are various for subjects. The generated questions should be customized to students Incorporate a knowledge tracking model [43, 44] with pre-defined profiles of students. Distractor Generation The produced distractors are irrelevant to the questions Employ dual attention and contrastive learning to ensure the consistency and high relevance to the contexts [45, 46, 47, 48, 49]. . The distractors lack strong distractive effectiveness Propose methods to avoid excessive similarity between the generated distractors [50, 51, 52]. . Automated Essay Scoring The singular evaluation of the essays lacks comprehensiveness Introduce more diverse evaluation metrics [53, 54] and explainable prompts [55] considering specific objects and scenarios. The sparse annotation of the essays in domain-specific prompts Decouple the domain-dependent and domain-independent features [56, 57, 58]. Automated Code Scoring It is hard to identify good features and scoring algorithms Collect comments/ number of lines/ number of operators to train a model to evaluate the quality of the code [59, 60, 61, 62, 63] with supervision . Implicitly encode the code snippet via distributed representation Encode a code snippet and compare with the correct ones [64, 65, 66, 67, 68, 69, 70] . Grammatical Error Correction GEC with LLMs encounters over-correction issue Conduct ensemble of the edits produced via LLMs [71, 72, 73]. Distorted results of GEC systems in low resources Construct multilingual data to handle mix-language text [74]. Code Error Correction The structural information of code has been ignored Utilize tree structure [75, 76, 77]/ AST data [wang-etal-2021-codet5] to model the structural characteristics of code. CEC systems have resource overhead issue Incorporating grammar constraints of programming languages [78, 79]. CEC systems encounter overfitting issue Evaluate on a wider range of execution scenarios [80]; propose overfitting detection systems [81, 82]
@@ 2 Taxonomy of NLP in Education Domain
Systems built on the basis of NLP applications and equipped with NLP techniques could perform as empowered AI tutoring, which enhances the teaching and learning experience of students and teachers. With the systems, students are able to get access to private tutoring and extra support outside the classroom, which avoids increasing the workload of teachers.
We first demonstrate an overview of NLP for education in Figure 1, where the systematic frame of relevant NLP techniques and their corresponding outcomes are displayed. The data sources are significant as the fundamental supports for subsequent analyses and applications. Generally, we could obtain the data from textbooks, syllabus, exams, and curricula materials, which may contain behavioral data of students, exercise questions, guidance signals of error correction, and so on. From the perspective of teaching subjects, the data resources cover subjects like chemistry, math, linguistics, computer science, and so on, which could support applications in different subjects. Given the raw data resources, we could conduct processing and analysis (e.g., knowledge graph construction, information extraction, information retrieval, knowledge alignment) to extract the key information from massive data and organize them in structured knowledge graphs (e.g., EduKG [^1] https://edukg.cn/team), databases (e.g., WordNet [^2] https://wordnet.princeton.edu/documentation) and lists, which could be applied to various downstream applications. Upon the well-organized data, NLP tasks are applied (e.g., question answering, question construction, automated assessment, and error correction) to form high-level applications to serve teachers and students in real-world scenarios. Grammarly [^3] https://app.grammarly.com, Gauthmath [^4] https://www.gauthmath.com, CodeSubmit [^5] https://codesubmit.io, AutoQG  [^6] https://autoqg.net, and EduChat [^7] https://www.educhat.top/ are possible educational products and demonstrations developed based on these NLP applications.
This survey mainly focuses on the application level. We highlight the most representative NLP applications with the taxonomy shown in Figure 3. Considering that NLP techniques play significant roles in different phases of education, on the one hand, NLP could serve as a teaching assistant for learning and comprehension. When students learn the curriculum, the following NLP tasks could simplify the difficulties of learning.
  • Question Answering (QA), as a typical application, could not only provide tutorials for students but also save time for overloaded teachers by providing feedback on questions. Based on the different goals of question answering, textbook question answering and math word problem solving are presented as the two typical subcategories of question answering. These two tasks both aim to generate answers from the given contexts automatically. They can be widespread in the education of science subjects.
  • Question Construction (QC) is to generate questions as well as the distractors given the correct answers, which helps the teachers create a wide range of test questions using pre-existing learning resources. Generally, a good test question may consist of a question and multiple distractors. Therefore, question generation and distractor construction are two key sub-tasks to be reviewed.
On the other hand, NLP improves writing and assessment for students. When given constructive criticism and writing prompts, students are likely to go deep into the topic. However, teachers sometimes have difficulties responding thoughtfully to students due to time limitations. But they will benefit from the following tasks.
@@
  • Error Correction (EC) for language are useful to improve language learning via explicating the correct edits. Still, taking language learning as an example, grammatical error correction and code error correction are canonical sub-tasks for EC. The methods are not simply required to provide error-free outputs based on the erroneous inputs, more detailed instructions should be provided to shed light on the reasons or knowledge points to the errors.
@@
Please be aware that there are some other NLP tasks that are related to education. For example, tasks like text rewriting [83] and knowledge graph construction [84] can also help teachers to prepare suitable teaching materials. In this survey, we mainly focus on the highlighted tasks, which are widely explored NLP tasks in the education domain. Our discussion is carried out based on the above taxonomy.